# tSAPPHIRES collaboration

### v1

@@ I Introduction
Also, phenomenologically there are indications favoring eV-scale ALPs. Recent measurements of the cosmic optical and infrared background (COB/CIB) have suggested possible excesses of diffuse emission in some bands. One intriguing interpretation attributes this excesses to two-photon decays of ALPs[11, 12, 13]. Independently, measurements of optical-depth of $\gamma$-rays from blazars show slightly stronger attenuation than that predicted by galaxy-count-based EBL models. Modeling the residual optical depth as photons from ALP two-photon decays favor eV-scale ALPs [14]. In the eV-mass region, several indirect-detection searches have been carried out [15, 16, 17, 18]. Among them, WINERED experiment used high-resolution spectroscopy to probe the $1.8~\text{-}~2.7~\mathrm{eV}$ window, setting stringent limits, and while the observing conditions were not ideal, several local $5\sigma$ features were noted [18].
@@
In this paper, as a laboratory-based search in the eV-mass scale, we present the result of the search based on stimulated resonant photon collisions (SRPC) [19, 20, 21, 22, 23] using three-laser beams. This method enables model-independent testing, free from astrophysical, astronomical and cosmological assumptions, by colliding two focused creation laser beams with a third focused inducing laser beam. The pair of creation lasers generates ALP-resonance states at the collision point. With the average energy of a single photon in the creation lasers, denoted by $\omega_{c}$, and subtended angles $2\theta_{c}$ between colliding two photons, the center-of-mass energy is given by $E_{cms}=2\omega_{c}sin\theta_{c}$. The inducing laser stimulates the decay of the generated ALP. Denoting the average single-photon energy in the inducing field as $\omega_{i}$, one of the two decay photons can be extracted as the signal photon with energy $\omega_{s}=2\omega_{c}-\omega_{i}$, which is kinematically equivalent to the atomic four-wave-mixing (FWM) process. From the center-of-mass energy formula, it follows that the ALP mass range can be changed by varying the single-photon energy $\omega_{c}$ and the incidence angle $\theta_{c}$. We have therefore proposed [24], peformed the pilot search [25] and developed a variable-incidence three-beam collider that enables a continuous eV-scale scan [26] (see also sub-eV searches in quasi-collinear, coaxially focused two-beam configurations performed by the SAPPHIRES Collaboration [27, 28, 29]).
@@ III Measurement
Figure 4 shows the arrival time distribution of FWM-like photons $N_{S}$ for the S pattern (500 shots) measured using PMT1 with NLC. The number of photons was reconstructed from the voltage-time relation of analog signals from PMT1 with a waveform digitizer. We also applied a peak-finding algorithm to simultaneously determine the number of photons and their arrival times from falling edges of amplitudes in the waveforms. The details of the peak-finding algolism are given in the Sec. IV. The arrival time of the FWM photons originating from the atomic process in NLC was used to define the time window corresponding to that for signal photon emission via ALP exchanges. With the creation and inducing lasers operated at low energy, the yields $N_{C}$ and $N_{I}$ in the C and I patterns were negligible. In the search experiment, we evaluated the number of photons within this red time window for counting the number of ALP-induced signal photons.
@@ IV Data analysis
Figure 5: Typical waveform and peak-finding algorithm measured by the digitizer. The orange dashed horizontal line shows the baseline mean $V_{0}$, estimated over the initial baseline window $t\in[0,100\,\mathrm{ns}]$. The red dashed horizontal line is the photon-detection threshold $L_{5\sigma}$. The green vertical line marks the peak time $t_{\mathrm{pk}}$ (time of the most negative sample that exceeds $L_{5\sigma}$). The cyan and blue dashed horizontal lines are the arrival-time threshold $L_{3\sigma}$ and the integration-start threshold $L_{2\sigma}$, respectively; these define the arrival time $t_{\mathrm{arr}}$ (cyan vertical line) and the integration-start time $t_{\mathrm{s}}$ (blue vertical line), respectively. The purple vertical line indicates the integration-end time $t_{\mathrm{e}}$, set to 3.5 ns after the peak time. The waveform charge is obtained by integrating from $t_{\mathrm{s}}$ to $t_{\mathrm{e}}$, then converted to photons using one-photon-equivalent charge (1-PEQ) from the PMT calibration (details in Appendix A). In this example, the estimated photon count is $\sim$ $1.$
@@
    Charge assignment. Compute the baseline-subtracted time integral between $t_{s}$ and $t_{e}$ with $\Delta t=0.5\,~\mathrm{ns}$ and $R=50~\Omega$,
@@
    and define this charge with the arrival timestamp $t_{\mathrm{arr}}$.
@@
    Photon conversion. Using the PMT single-photoelectron calibration (Appendix A), we define the one-photon-equivalent charge $Q_{1\mathrm{PEQ}}$. The corresponding number of detected photons is
    $N\;=\;\frac{Q}{Q_{1\mathrm{PEQ}}}\,.$ (8)
@@
$n_{\mathrm{obs}}=-30.5\pm 45.0\,\text{(stat.)}\pm 13.9\,\text{(syst.\,I)}$ (9)
@@
where the three systematic errors are coming from the definition of the peak-finding algorithm as follows. The first systematic error (syst.I) was evaluated by varying the voltage threshold level in Eq. (3) used to discriminate photon-peaks from baseline fluctuations,
@@ V Upper limit in the coupling-mass relation for ALP-exchange
However, experimentally, the overlap factor $\mathcal{D}_{three}$ is expected to deteriorate due to variations in the incidence angle of each beam, $\Delta{\theta_{x,j}},\Delta{\theta_{y,j}},(j=c1,c2,i)$, which in turn causes shifts in peak positions at the focal point, $\Delta{x}_{j},\Delta{y}_{j},c\Delta{t}_{j},(j=c1,c2,i)$. Therefore, in the search, the peak positions of the focal spot profiles of the three laser beams were measured before and after each run r = 1,2,3 using IPM2. The focal position measured before each of the three runs was treated as the ideal overlap factor with zero misalignment, denoted as $\mathcal{D}_{0}$. The overlap factor after each run including the three beam offsets is expressed using the measured parameters as $\mathcal{D}_{r}(\Delta{x}_{j},\Delta{y}_{j},c\Delta{t}_{j},\Delta{\theta_{x,j}},\Delta{\theta_{y,j}})$. Here, the reduction in the overlap factor from timing and angular mismatches $c\Delta{t}_{j},\Delta{\theta_{x,j}},\Delta{\theta_{y,j}}$ is much smaller than that from the transverse offsets $\Delta{x}_{j},\Delta{y}_{j}$. We therefore retain only the $\Delta x_{j},\ \Delta y_{j}$ contributions. Finally, the merged and averaged overlap factor $\overline{\mathcal{D}}_{\mathrm{three}}$, is defined from the pre- and post-run values as
@@
with $N_{\text{run}}=3$ in this experiment. Typical transverse shifts were $\Delta x_{j},\Delta y_{j}\simeq\pm 3\,\mu\mathrm{m}$ for $j\in\{c1,c2,i\}$. For reference, the measured waist diameters at the CP were $2w_{c}\simeq 30\,\mu\mathrm{m}$ (creation beams) and $2w_{i}\simeq 52\,\mu\mathrm{m}$ (inducing beam). Further details of the three-beam overlap factor, $\overline{\mathcal{D}}_{\text{three}}$, are given in Appendix B. Therefore, the experimental yield $\mathcal{Y}_{c+i}$ in Eq. (13) is re-expressed as
@@
Duration time of inducing laser beam (FWHM), $\sqrt{2\ln 2}~\tau_{ibeam}$ 7.11 ns
@@
$\delta n_{obs}=\sqrt{45.0^{2}+13.9^{2}+25.5^{2}+30.6^{2}}\simeq 62.$ (18)
@@
Figure 8: Top: ALP-favored regions from COB/CIB and $\gamma$-ray attenuation are shown; the gray region shows blazar $\gamma$-ray attenuation (ALP-decay model) [14], the cyan and yellow bands show COB excess (New Horizons/LORRI ’23,’24) [13], orange region shows near-IR anisotropies (HST+Spitzer; ALP+IHL fit) [12]. The beige band and the upper solid brown line are the predictions of QCD axion by the KSVZ model [33] with $0.07<\left|E/N-1.95\right|<7$ and $E/N=0$, respectively. The bottom solid brown line is the prediction from the DFSZ model [34] with $E/N=8/3$. The cyan lines are the predictions from the ALP miracle model [8] with the intrinsic parameter values $c_{\gamma}=1,0.1,0.01$, respectively. Bottom: Upper limit in the parameter space of the coupling-mass relation (region enclosed by the red solid curve) evaluated at a 95% confidence level for the pseudoscalar field exchange shown as tSAPPHIRES01. The previous search, tSAPPHIRES00, presents the excluded region based on the runs conducted in air at $\theta_{c}=30^{\circ}$ by the red solid curve [25]. The magenta area shows the excluded range based on two-beam SRPC in quasi-parallel collision geometry (SAPPHIRES02) [29]. The purple areas are excluded regions by the Light-Shining-through a Wall (LSW) experiments (ALPS [35] and OSQAR [36]). The gray area shows the excluded region by the vacuum magnetic birefringence (VMB) experiment (PVLAS [37]). The light-cyan horizontal solid line indicates the upper limit from the search for eV pseudoscalar penetrating particles in the SPS neutrino beam (NOMAD) [38]. The green area is the excluded region by the helioscope experiment CAST [39]. The horizontal dotted line indicates the upper limit from the Horizontal Branch observation [40, 41]. the light-green regions indicate exclusions from JWST/NIRSpec adopting NFW profile of a Milky Way DM halo [15]. The purple region indicates exclusion by WINERED spectroscopy adopting NFW profile of dSphs (Leo V, Tucana II); localized $5\sigma$ excesses remain unexplained [18]. The blue region indicates exclusion by VLT/MUSE-Faint observations adopting NFW profile of Leo T [16].
@@
Figure 8 (top) shows the ALP-favored regions and theoretical models: the gray region shows the parameter space favored by $\gamma$-ray attenuation of blazar spectra when the optical-depth excess is modeled as ALP two-photon decays [14]. The cyan and yellow regions show the parameter spaces favored by COB excess (New Horizons/LORRI ’23,’24) [13]. The orange regions show the parameter spaces favored by HST + Spitzer near-IR anisotropy data when fit with an ALP + intra-halo-light [12]. The beige band is the region of the benchmark QCD-axion model (KSVZ [31, 32]) with $0.07<\left|E/N-1.95\right|<7$. The upper solid brown line denotes the KSVZ model [33] with $E/N=0$, whereas the lower solid brown line denotes the DFSZ model [34] with $E/N=8/3$. The cyan lines are the predictions from the ALP miracle model [8] with the intrinsic parameters $c_{\gamma}=1,0.1,0.01$.
@@
Figure 8 (bottom) shows the upper limit in the coupling-mass plane from this search, tSAPPHIRES01, enclosed by the red solid curve. The limit was set at a 95% confidence level for pseudoscalar-type ALP exchanges. The peak sensitivity was at $m_{a}=2.27~\mathrm{eV}$ with $g/M\sim 4.0\times 10^{-10}\,\mathrm{GeV}^{-1}$, using a central wavelength $\lambda_{c}=811~\mathrm{nm}$ and an incidence angle $\theta_{c}=47.9^{\circ}$. This represents the first model-independent exclusion/sensitivity reaching KSVZ line ($E/N=0$) [33]. The mass broadening at a relaxed coupling of $g/M\sim 4.0\times 10^{-9}\,\mathrm{GeV}^{-1}$ was approximately $m_{a}=2.27\pm 0.07~\mathrm{eV}$. The sensitivity spread arises from fluctuations in the center-of-mass energy, driven by the finite momentum spread and bandwidth of the focused, short-pulse lasers [24]. The previous search tSAPPHIRES00 presents the excluded region based the runs conducted in air at an incidence angle $\theta_{c}=30^{\circ}$ (corresponding to $m_{a}=1.54,\mathrm{eV}$) [25]. We note that the label tSRPC00 used in Ref.[25] has been relabeled as tSAPPHIRES00 following reanalysis, to ensure consistency with tSAPPHIRES01 in the derivation of experimental parameters and the beam overlap factor.
@@
The magenta area indicates the excluded range based on SRPC in quasi-parallel collision geometry (SAPPHIRES02) [29]. The purple areas are excluded regions by the Light-Shining-through a Wall (LSW) experiments (ALPS [35] and OSQAR [36]). The gray area shows the excluded region by the vacuum magnetic birefringence (VMB) experiment (PVLAS [37]). The light-cyan horizontal solid line indicates the upper limit from the search for eV pseudoscalar penetrating particles in the SPS neutrino beam (NOMAD) [38]. The green area is the excluded region by the helioscope experiment CAST [39]. The horizontal dotted line indicates the upper limit from the Horizontal Branch observation [40, 41]. The light-green regions indicate exclusions from JWST/NIRSpec-IFU blank-sky spectra [15]. These limits overlap our search region but depend on background modeling (smooth continuum model) and astrophysical assumptions about Milky Way DM halo (NFW), hence complementary. The purple regions indicate exclusions from high-dispersion WINERED spectroscopy of dSphs (Leo V, Tucana II adopting NFW profile) [18]. However, the authors also noted localized $5\sigma$ excesses in that window, and its origin remains uncertain. The blue regions indicate exclusions from the VLT/MUSE-Faint observations of a dSph (Leo T adopting NFW profile) [16].

### v2

@@ I Introduction
Also, phenomenologically there are indications favoring eV-scale ALPs. Recent measurements of the cosmic optical and infrared background (COB/CIB) have suggested possible excesses of diffuse emission in some bands. One intriguing interpretation attributes this excesses to two-photon decays of ALPs[11, 12, 13]. Independently, measurements of optical-depth of $\gamma$-rays from blazars show slightly stronger attenuation than that predicted by galaxy-count-based EBL models. Modeling the residual optical depth as photons from ALP two-photon decays favors eV-scale ALPs [14]. In the eV-mass region, several indirect-detection searches have been carried out [15, 16, 17, 18]. Among them, WINERED experiment used high-resolution spectroscopy to probe the $1.8~\text{-}~2.7~\mathrm{eV}$ window, setting stringent limits, and while the observing conditions were not ideal, several local $5\sigma$ features were noted [18].
@@
In this paper, as a laboratory-based search in the eV-mass scale, we present the result of the search based on stimulated resonant photon collisions (SRPC) [19, 20, 21, 22, 23] using three-laser beams. Figure 1 shows a schematic view of the three-beam stimulated resonant photon collider (tSRPC). This method enables model-independent testing, free from astrophysical, astronomical and cosmological assumptions, by colliding two focused creation laser beams with a third focused inducing laser beam. The pair of creation lasers generates ALP-resonance states at the collision point. With the average energy of a single photon in the creation lasers, denoted by $\omega_{c}$, and subtended angles $2\theta_{c}$ between colliding two photons, the center-of-mass system energy is given by
$E_{cms}=2\omega_{c}\sin\theta_{c}.$ (1)
The inducing laser stimulates the decay of the generated ALP. Denoting the average single-photon energy in the inducing field as $\omega_{i}$, one of the two decay photons can be extracted as the signal photon with energy $\omega_{s}=2\omega_{c}-\omega_{i}$, which is kinematically equivalent to the atomic four-wave-mixing (FWM) process. In practice, photons from each beam collide with a spread of angles and energies: focusing induces transverse-momentum (angular) spread, and the ultrashort pulse duration gives a finite energy bandwidth. As a result, the center-of-mass system energy is effectively modified to
$E_{cms}=2\sqrt{\omega_{1}\omega_{2}}\sin\left(\frac{\vartheta_{1}+\vartheta_{2}}{2}\right),$ (2)
and a fraction of the inducing field $\omega_{4}$ stimulates scattering into signal photons of energy $\omega_{3}$ that satisfy energy-momentum conservation. Consequently, the collision energy is not fixed but distributed, and by changing the incidence angle $\theta_{c}$ with a step matched to the mass resolution, we can perform a fast and continuous mass scan. We have therefore proposed [24], peformed the pilot search [25] and developed a variable-incidence three-beam collider that enables a continuous eV-scale scan [26] (see also sub-eV searches in quasi-collinear, coaxially focused two-beam configurations performed by the SAPPHIRES Collaboration [27, 28, 29]).
Figure 1: Concept of the three-beam stimulated resonant photon collider (tSRPC). Three short-pulsed laser beams with photon number densities $\rho_{k}(k=c_{1},c_{2},i)$ are independently focused at the collision point. The two creation beams (green) cross at an incidence angle $\theta_{c}$ and produce an ALP resonance, while the inducing beam (red) stimulates its decay. The creation photons have energies $\omega_{1}$ and $\omega_{2}$ and incidence angles $\theta_{1}$ and $\theta_{2}$ around the central values $\omega_{c}$ and $\theta_{c}$, respectively. Similarly, the inducing beam with central energy $\omega_{i}$ contains a frequency component $\omega_{4}$ that enhances emission of signal photons with energy $\omega_{3}$ (blue) via energy-momentum conservation.
@@ III Measurement
Figure 5 shows the arrival time distribution of FWM-like photons $N_{S}$ for the S pattern (500 shots) measured using PMT1 with NLC. The number of photons was reconstructed from the voltage-time relation of analog signals from PMT1 with a waveform digitizer. We also applied a peak-finding algorithm to simultaneously determine the number of photons and their arrival times from falling edges of amplitudes in the waveforms. The details of the peak-finding algorithm are given in the Sec. IV. The horizontal axis represents the photon arrival time with a bin width of $\Delta t=0.5$ ns, and the vertical axis shows the number of reconstructed photons per 0.5 ns. The arrival time of the FWM photons originating from the atomic process in NLC was used to define the time window corresponding to that for signal photon emission via ALP exchanges. With the creation and inducing lasers operated at low energy, the yields $N_{C}$ and $N_{I}$ in the C and I patterns were negligible. In the search experiment, we evaluated the number of photons within this red time window for counting the number of ALP-induced signal photons.
@@ IV Data analysis
Figure 6: Typical waveform and peak-finding algorithm measured by the digitizer. The orange dashed horizontal line shows the baseline mean $V_{0}$, estimated over the initial baseline window $t\in[0,100\,\mathrm{ns}]$. The red dashed horizontal line is the photon-detection threshold $L_{5\sigma}$. The green vertical line marks the peak time $t_{\mathrm{pk}}$ (time of the most negative sample that exceeds $L_{5\sigma}$). The cyan and blue dashed horizontal lines are the arrival-time threshold $L_{3\sigma}$ and the integration-start threshold $L_{2\sigma}$, respectively; these define the arrival time $t_{\mathrm{arr}}$ (cyan vertical line) and the integration-start time $t_{\mathrm{s}}$ (blue vertical line), respectively. The purple vertical line indicates the integration-end time $t_{\mathrm{e}}$, set to 3.5 ns after the peak time. The waveform charge is obtained by integrating from $t_{\mathrm{s}}$ to $t_{\mathrm{e}}$, then converted to photons using one-photon-equivalent charge from the PMT calibration (details in Appendix A). In this example, the estimated photon count is $\sim$ $1.$
@@
    Charge assignment. Compute the baseline-subtracted time-integral voltage between $t_{s}$ and $t_{e}$ with $\Delta t=0.5~\mathrm{ns}$ as
@@
    where $R=50~\Omega$ is the nominal input termination of the digitizer; since $R$ cancels in the calibration ratio, its precise value does not affect the final results. The resulting charge $Q$ is then assigned to the arrival timestamp $t_{\mathrm{arr}}$.
@@
    Photon conversion. Using the single-photoelectron calibration of the PMT (Appendix A), performed with the same PMT, cabling, and front-end electronics as in the search data, we determine the single-photon-equivalent charge to be
    $Q_{1\mathrm{PE}}=-0.2337\pm 0.0004~[\mathrm{pC/photon}].$ (10)
    The corresponding number of detected photons is
    $N\;=\;\frac{Q}{Q_{1\mathrm{PE}}}\,.$ (11)
@@
$n_{\mathrm{obs}}=-30.5\pm 52.4\,\text{(stat.)}\pm 13.9\,\text{(syst.\,I)}$ (12)
@@
where the first uncertainty is statistical, evaluated from the Poisson fluctuations of the reconstructed photon numbers in the four trigger patterns, including the Fano-factor enhancement due to the Gamma-distributed single-photon charge response, as given by Eq.(42) and discussed in Appendix B, while the three systematic errors are coming from the definition of the peak-finding algorithm as follows. The first systematic error (syst.I) was evaluated by varying the voltage threshold level in Eq. (5) used to discriminate photon-peaks from baseline fluctuations,
@@ V Upper limit in the coupling-mass relation for ALP-exchange
The center–of–mass system energy $E_{\mathrm{cms}}$ of two colliding laser photons is intrinsically smeared by quantum uncertainty: each photon must be treated as a wave packet with finite temporal duration, which implies a finite spectral width via the Fourier-transform relation between time and energy (set by the pulse width), and with a finite angular spread, which is Fourier-related to the beam waist (set by the focusing optics). In contrast, the natural ALP decay rate, $\Gamma=(g/M)^{2}m_{a}^{3}/(16\pi)$, that enters the Breit–Wigner propagator is extremely small in the coupling range of interest. For example, for $g/M=10^{-10}\,\mathrm{GeV}^{-1}$ and $m_{a}=1\,\mathrm{eV}$, one finds $\Gamma\simeq 2\times 10^{-40}\,\mathrm{eV}$. Compared to any realistic experimental window $\Delta E_{\mathrm{cms}}$ on $E_{\mathrm{cms}}$, the Breit-Wigner width is therefore almost delta-function-like. However, the integrated effect of the s-channel pole, including the off-shell part of the Breit-Wigner function within $\Delta E_{\mathrm{cms}}$, can drastically increase the sensitivity [23]. In our scheme, the observable lineshape is in practice governed by the experimental distribution of $E_{\mathrm{cms}}$, rather than by the tiny intrinsic ALP width. Given an accurately controlled $\Delta E_{\mathrm{cms}}$ in the experiment, if an ALP resonant mass lies within that range, we can claim the existence of a resonance through the averaged pole contribution.
Experimentally, the overlap factor $\mathcal{D}_{three}$ is expected to deteriorate due to variations in the incidence angle of each beam, $\Delta{\theta_{x,j}},\Delta{\theta_{y,j}},(j=c1,c2,i)$, which in turn causes shifts in peak positions at the focal point, $\Delta{x}_{j},\Delta{y}_{j},c\Delta{t}_{j},(j=c1,c2,i)$. Therefore, in the search, the peak positions of the focal spot profiles of the three laser beams were measured before and after each run r = 1,2,3 using IPM2. The focal position measured before each of the three runs was treated as the ideal overlap factor with zero misalignment, denoted as $\mathcal{D}_{0}$. The overlap factor after each run including the three beam offsets is expressed using the measured parameters as $\mathcal{D}_{r}(\Delta{x}_{j},\Delta{y}_{j},c\Delta{t}_{j},\Delta{\theta_{x,j}},\Delta{\theta_{y,j}})$. Here, the reduction in the overlap factor from timing and angular mismatches $c\Delta{t}_{j},\Delta{\theta_{x,j}},\Delta{\theta_{y,j}}$ is much smaller than that from the transverse offsets $\Delta{x}_{j},\Delta{y}_{j}$. We therefore retain only the $\Delta x_{j},\ \Delta y_{j}$ contributions. Finally, the merged and averaged overlap factor $\overline{\mathcal{D}}_{\mathrm{three}}$, is defined from the pre- and post-run values as
@@
with $N_{\text{run}}=3$ in this experiment. Typical transverse shifts were $\Delta x_{j},\Delta y_{j}\simeq\pm 3\,\mu\mathrm{m}$ for $j\in\{c1,c2,i\}$. For reference, the measured waist diameters at the CP were $2w_{c}\simeq 30\,\mu\mathrm{m}$ (creation beams) and $2w_{i}\simeq 52\,\mu\mathrm{m}$ (inducing beam). Further details of the three-beam overlap factor, $\overline{\mathcal{D}}_{\text{three}}$, are given in Appendix C. Therefore, the experimental yield $\mathcal{Y}_{c+i}$ in Eq. (16) is re-expressed as
@@
Duration time of inducing laser beam (FWHM), $\sqrt{2\ln 2}~\tau_{ibeam}$ 7.1 ns
@@
$\delta n_{obs}=\sqrt{52.4^{2}+13.9^{2}+25.5^{2}+30.6^{2}}\simeq 67.$ (21)
@@
Figure 9: Top: ALP-favored regions from COB/CIB and $\gamma$-ray attenuation are shown; the gray region shows blazar $\gamma$-ray attenuation (ALP-decay model) [14], the cyan and yellow bands show COB excess (New Horizons/LORRI ’23,’24) [13], orange region shows near-IR anisotropies (HST+Spitzer; ALP+IHL fit) [12]. The beige band and the upper brown solid line correspond to QCD-axion predictions in the KSVZ framework [33], for $0.07<\left|E/N-1.95\right|<7$ and $E/N=0$, respectively. The lower brown solid line shows the DFSZ prediction [34] with $E/N=8/3$. The cyan lines indicate the ALP miracle model [8] for $c_{\gamma}=1,0.1,0.01$. Bottom: The red solid contour encloses the 95% confidence level upper-limit region in the coupling-mass plane for pseudoscalar exchange (labelled tSAPPHIRES01). The previous search, tSAPPHIRES00, presents the excluded region based on the runs conducted in air at $\theta_{c}=30^{\circ}$ by the red solid curve [25]. The magenta region shows the SRPC exclusion in a quasi-parallel two-beam configuration (SAPPHIRES02) [29] The purple shaded regions show the excluded parameter space from Light-Shining-through-a-Wall (LSW) searches by ALPS [35] and OSQAR [36]. The gray region indicates the limit from the vacuum magnetic birefringence (VMB) measurement by PVLAS [37]. The light-cyan horizontal solid line denotes the NOMAD upper limit from the SPS neutrino beam search [38]. The green region shows the exclusion from the CAST helioscope [39]. The horizontal dotted line denotes the upper bound from the HB observation [40, 41]. The light-green regions indicate exclusions from JWST/NIRSpec adopting NFW profile of a Milky Way DM halo [15]. The purple region indicates exclusion by WINERED spectroscopy adopting NFW profile of dSphs (Leo V, Tucana II); localized $5\sigma$ excesses remain unexplained [18]. The blue region indicates exclusion by VLT/MUSE-Faint observations adopting NFW profile of a sample of five galaxies [16].
@@
Figure 9 (top) shows the ALP-favored regions and theoretical models: the gray region shows the parameter space favored by $\gamma$-ray attenuation of blazar spectra when the optical-depth excess is modeled as ALP two-photon decays [14]. The cyan and yellow regions show the parameter spaces favored by COB excess (New Horizons/LORRI ’23,’24) [13]. The orange regions show the parameter spaces favored by HST + Spitzer near-IR anisotropy data when fit with an ALP + intra-halo-light [12]. The beige shaded band corresponds to the benchmark QCD-axion model (KSVZ [31, 32]) with $0.07<\left|E/N-1.95\right|<7$. The upper brown solid line shows the KSVZ model [33] with $E/N=0$, while the lower brown solid curve indicates the DFSZ model [34] with $E/N=8/3$. The cyan lines illustrate the predictions of the ALP miracle scenario [8] with the intrinsic parameters $c_{\gamma}=1,0.1,0.01$.
@@
Figure 9 (bottom) shows the upper limit in the coupling-mass plane from this search, tSAPPHIRES01, enclosed by the red solid curve. The limit was set at a 95% confidence level for pseudoscalar-type ALP exchanges. The peak sensitivity was at $m_{a}=2.27~\mathrm{eV}$ with $g/M\sim 4.2\times 10^{-10}\,\mathrm{GeV}^{-1}$, using a central wavelength $\lambda_{c}=811~\mathrm{nm}$ and an incidence angle $\theta_{c}=47.9^{\circ}$. This represents the first model-independent exclusion/sensitivity reaching KSVZ line ($E/N=0$) [33]. The mass broadening at a relaxed coupling of $g/M\sim 4.2\times 10^{-9}\,\mathrm{GeV}^{-1}$ was approximately $m_{a}=2.27\pm 0.07~\mathrm{eV}$. The sensitivity spread arises from fluctuations in the center-of-mass system energy, driven by the finite momentum spread and bandwidth of the focused, short-pulse lasers [24]. The previous search tSAPPHIRES00 presents the excluded region based the runs conducted in air at an incidence angle $\theta_{c}=30^{\circ}$ (corresponding to $m_{a}=1.54,\mathrm{eV}$) [25]. We note that the label tSRPC00 used in Ref.[25] has been relabeled as tSAPPHIRES00 following reanalysis, to ensure consistency with tSAPPHIRES01 in the derivation of experimental parameters and the beam overlap factor.
@@
The magenta shading shows the parameter space excluded by SRPC searches performed in a quasi-parallel collision setup (SAPPHIRES02) [29]. The purple shaded regions correspond to constraints from Light-Shining-through-a-Wall (LSW) experiments, including results from ALPS [35] and OSQAR [36]. The gray region indicates the exclusion from the vacuum magnetic birefringence (VMB) measurement (PVLAS [37]). The light-cyan horizontal solid line shows the upper bound obtained by NOMAD using the SPS neutrino beam [38]. The green shaded region denotes the CAST helioscope exclusion [39], obtained from searches for solar axions/ALPs via axion-photon conversion. The horizontal dotted line is disfavoured by globular-cluster constraints based on the impact of ALP emission on stellar evolution [40, 41]. The light-green regions indicate exclusions from JWST/NIRSpec-IFU blank-sky spectra [15]. These limits overlap our search region but depend on background modeling (smooth continuum model) and astrophysical assumptions about Milky Way DM halo (NFW), hence complementary. The purple regions indicate exclusions from high-dispersion WINERED spectroscopy of dSphs (Leo V, Tucana II adopting NFW profile) [18]. However, the authors also noted localized $5\sigma$ excesses in that window, and its origin remains uncertain. The blue regions indicate exclusions from the VLT/MUSE-Faint observations of a dSph (Leo T adopting NFW profile) [16].